Given the following macro-economics model :

Y = C+I+G

C = 20+0.07Y

I = 12+0.1Y

G = 10


a. Express the model in matrix form.

b. Using Cramer's Rule , calculate the equilibrium values of national income (Y) , consumption (C) and investment (I)

c. what is the income multiplier?


1
Expert's answer
2021-04-06T07:27:23-0400

Y = C+I+G

C = 20+0.07Y

I = 12+0.1Y

G = 10


(a) In matrix form Itcan be represented as-


"\begin{pmatrix}
 1 & -1&-1&0 \\
 -0.07&1&0&0 \\
 0&0&1&-0.1\\
 K&0&0&-l
\end{pmatrix}
\begin{pmatrix}
 Y \\C \\
 I\\ i
\end{pmatrix}=\begin{pmatrix}
 G_o\\20 \\
 12 \\ M_o
\end{pmatrix}"


(b) As National Income


"Y=C+I+G\\

 Y =20+0.07Y+12+0.1Y+10\\

 Y =42+0.17Y"

"0.83Y=42\Rightarrow Y=50.60"


Consumption, "C=20+0.07(50.60)=23.542"


Investment, "I=12+0.1(50.60)=17.06"


(c) Income Multiplier "=\dfrac{C}{I}=\dfrac{23.542}{17.06}=1.38"